OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. a14 = -33, a15 = 9 an = -579 + 42(n + 1) an = -579 + 42(n - 1) an = -579 - 42(n + 1) an = -579 - 42(n - 1)

OpenStudy (cwrw238):

common difference is a15 - a14

OpenStudy (cwrw238):

nth term = a1 +( n-1)d

OpenStudy (cwrw238):

common difference d = 9 - (-33)

OpenStudy (cwrw238):

can you calculate d = 9-(-33)?

OpenStudy (anonymous):

42?

OpenStudy (anonymous):

so the answer would be an = -579 + 42(n - 1) ??? is this right?

OpenStudy (cwrw238):

yes